### v1

# Why do explanations fail? A typology and discussion on failures in XAI

Abstract
As Machine Learning (ML) models achieve unprecedented levels of performance, the XAI domain aims at making these models understandable by presenting end-users with intelligible explanations. Yet, some existing XAI approaches fail to meet expectations: several issues have been reported in the literature, generally pointing out either technical limitations or misinterpretations by users. In this paper, we argue that the resulting harms arise from a complex overlap of multiple failures in XAI, which existing ad-hoc studies fail to capture. This work therefore advocates for a holistic perspective, presenting a systematic investigation of limitations of current XAI methods and their impact on the interpretation of explanations. By distinguishing between system-specific and user-specific failures, we propose a typological framework that helps revealing the nuanced complexities of explanation failures. Leveraging this typology, we also discuss some research directions to help AI practitioners better understand the limitations of XAI systems and enhance the quality of ML explanations.
Introduction
The field of eXplainable Artificial Intelligence (XAI) aims at addressing the challenge of providing users with explanations regarding decisions of Machine Learning (ML) models, bridging the gap between the inner workings of complex algorithms and human understanding. It constitutes a multidisciplinary domain, drawing upon not only computer science but also cognitive sciences, philosophy and human-computer interaction (Miller 2019).
Central to the study of XAI is the elusive concept of a ”good” explanation. Formal definitions (Amgoud and Ben-Naim 2022) do not capture the human component and prove to be a daunting task, leading researchers to gradually develop ad-hoc desiderata and investigate specific challenges that arise. The domain multidisciplinarity often results in fragmented investigations, with different research communities focusing on disparate aspects of the problem. On one side, some limitations of explainers such as their lack of robustness (Alvarez-Melis and Jaakkola 2018a) or faithfulness (Jacovi and Goldberg 2020) have been mostly investigated by the ML community. In parallel, other works in different fields (e.g. social sciences) have pointed out issues such as the difficulty for explanations to meet user needs (Matarese, Rea, and Sciutti 2021), prior beliefs (Riveiro and Thill 2022) or general reasoning (Bertrand et al. 2022).
However, this fragmented approach poses a significant challenge, as the limitations that XAI systems face are not mutually exclusive; we argue that they may overlap and conflate with each other. Consequently, AI practitioners may find themselves unable to comprehend the origin of a failure, and thus to properly mitigate the resulting harms. To circumvent this issue, we argue that a holistic approach to XAI failures is needed. To do so, we rely on a literature review of recent papers on XAI approaches, interfaces and evaluations, that identify in an ad hoc manner some limitations to the intelligibility of ML explanations. We propose a typology that provides a systematic and hierarchically organized overview of XAI failures, distinguishing between system-specific and user-specific ones. We believe that this typology can help AI practitioners to have a deeper understanding of their connections and origins. Leveraging this typology, we discuss research avenues to enhance the quality of explanations provided by XAI systems. By fostering a more nuanced understanding of the limitations inherent to XAI, we hope to pave the way for more effective and transparent automated decision-making processes.
This paper is organised as follows: the first section discusses the context and motivation for a holistic view of XAI failures, and provides an overview of the typology of XAI failures we propose. We then describe its two main categories in the two following sections respectively. Finally, we discuss in the last section avenues for future research.
Context and objective
Throughout the paper, we consider that the explanation process is composed of two components of different nature: (i) the ML system and (ii) the users.
Machine Learning System
On one side there is the ML system composed by a ML model providing decisions and an explanation method generating rationale for these predictions. We consider a supervised learning context, where given some input information, a ML model returns an associated decision. The data used as input can be either structured (e.g. tabular) or unstructured (images, text, etc.). The nature of the model can vary a lot, from simple (e.g. linear) models to deep neural networks and large language models. The model performance is evaluated using some metrics of performance; in the paper we focus on accuracy. In addition to the prediction itself, the system provides rationale for it, through an explanation generator. It can either be the predictive model itself (using transparent models) or a separate system, called post-hoc, built on top of the predictive model. There exists a huge diversity of methods to generate various kinds of information that act as explanations (see e.g. (Dwivedi et al. 2023) for a recent survey), either for one prediction (local explanation) or for the whole model behavior (global explanation). Additionally, some recent works propose to design explanation user interfaces (XUIs) (Chromik and Butz 2021), to display the generated explanation to the end-users in an intelligible and useful manner.
The users
The counterpart of this two-part explanation process is the user who receives the explanation and interacts with the ML system to accomplish their task. The explanation should allow them to understand the decisions made by the model. It must be underlined that users can have various objectives, expertise levels and prior knowledge (Liao, Gruen, and Miller 2020), defining their needs in terms of interpretability (Mohseni, Zarei, and Ragan 2018).
In this explanation process, we consider that a successful explanation thus depends on three elements: the ability (i) of the model to make an accurate prediction, (ii) of the explainer to provide a faithful explanation that addresses user needs, and (iii) of the user to properly understand and use it. When at least one of these elements fails, we say that there is an explanation failure that needs to be investigated. Most of the literature about XAI limitations, detailed in the next sections, has focused on discussing such failures by adopting a technical point of view (see e.g. (Molnar et al. 2020) for an overview). Very few works suggest that limitations coming from the user side should also be investigated, such as mismatches between explanations and user needs (Matarese, Rea, and Sciutti 2021) or biased inferences (Bertrand et al. 2022).
In addition, most existing works consist in ad hoc studies on specific problems, viewing them as independent from one another. Yet, given the interactive nature of the explanation process (Miller 2019), it is likely that addressing issues in a more global manner is needed: the process of an explanation is sequential, with first the ML system providing predictions and explanations and then the user interpreting and possibly interacting with them. It can be expected that some failures may interact, conflate, or even amplify one another.
Based on a review of recent literature in XAI, we propose to structure explanation failures in a three-level hierarchical structure, graphically represented in Fig. 1 and 2 and discussed in the next sections. At the first level, we propose to classify explanation failures into two categories: the system-specific one, discussed in the next section, groups issues associated to technical limitations of the ML system; the user-specific one, discussed in the following section, encompasses issues caused by the inferences users make about the provided explanations, that are specific to each user. In each case we provide information on the potential effects with examples from the literature, only referring to few references due to place constraints.
Figure 1: Types of system-specific explanation failures: misleading explanations ascribed to ML system and contradicting ones ascribed to the explainer. The cross and the $\neq$ symbol indicate the step at which the failure takes place.
System-specific failures
We identify system-specific explanation failures as a design problem of the ML system, that may be ascribed to all of its components, as represented on Fig. 1. We propose to structure them into two types of system-specific failures: (i) misleading explanations when the ML model is providing an inaccurate prediction or when the explainer is not faithful; and (ii) contradictions when conflicting information are provided by one or several explainers.
Misleading explanations
We call an explanation misleading when the failure results from the ML system being dysfunctional, i.e. when it fails to meet the very purpose it was designed for. We distinguish two situations, depending on whether the dysfunction comes from the prediction or the explanation. There is a risk that the explanation, however, could be accepted by the users without them being able to perceive this failure (”I understand — but should I?” in Fig. 1).
Why does it happen?
First, it can occur that ML models output confident yet incorrect predictions. In this case we say the ML model is dysfunctional, raising the question of the relevance or potential misleadingness of generating explanations. Second, the XAI system is deemed dysfunctional, when it fails to meet the mathematical objectives it has been designed to satisfy: most explanations are generated to satisfy required properties pre-defined by the ML practitioner (e.g., optimize certain objectives). Some of them, such as gradient-based methods (e.g., (Selvaraju et al. 2017)), are directly defined as closed-form formulas. In this case, we consider that they are functional by design, and the resulting explanations faithful. Hence, the issues these methods may encounter fall into other types of failures, detailed later. On the other hand, some explanation methods rely on optimizing cost functions (see e.g., counterfactual example generation (Mothilal, Sharma, and Tan 2020)), and as such do not guarantee that the associated desiderata is satisfied. Often, there is no guarantee that a satisfying solution to these formalized problems exists, as discussed for instance by (Ye and Durrett 2022) for prompt-based explanations for Large Language Models. As a result, the explanation may be unfaithful to the model it aims at explaining.
Why is it a problem?
When an incorrect prediction is returned, although explanations can be useful for model calibration (Ye and Durrett 2022), they may be seen as potentially harmful, especially for users with low levels of awareness (Papenmeier, Englebienne, and Seifert 2019). An unfaithful explanation is also obviously problematic, as a lack of fidelity to the model may induce the user either to reject the system or to place unwarranted trust in it (Papenmeier, Englebienne, and Seifert 2019).Yet, despite its critical importance, this faithfulness is often overlooked by both users and ML practitioners (Kaur et al. 2020). Even when there is a will to control for faithfulness, the precise definition and evaluation of this notion remains elusive (Jacovi and Goldberg 2020), amplifying the challenge. We argue that assessing faithfulness is crucial, and that this evaluation should precede all other assessments of the explanation: firstly, contrary to other assessments, it is a purely technical task, allowing ML developers to conduct it independently of end-users. Secondly, it serves as a foundational step for identifying and addressing any other potential issue. In the rest of the paper, the ML system is therefore assumed to be functional, i.e. to provide accurate predictions and faithful explanations.
Contradictory explanations
The explanation failures that we group under the term contradictory explanations are characterized by conflicting information provided by one or several explainers, that are assumed to provide faithful information with respect to an accurate ML model to explain. More precisely, these inconsistencies occur when there is a contradiction between different pieces of information of the same explanation (competing explanations), between different explanations by the same explainer (unstable explanations), or different explanations by different explainers (incompatible explanations). One consequence of this limitation is that users may not understand this conflicting information, and wonder ”Why is it different?” (bottom part of Fig 1).
Competing explanations
We say that explanations are competing when several parts of the explanation are contradicting with one another.
Why does it happen?
We identify two scenarios when this can happen. Often, explanations are composed of several components, interacting with each other in various ways: e.g., a feature attribution vector represents the contributions of each feature, word or pixel to the prediction. Yet, numerous works show that more complex effects such as interactions or correlations between features are often not taken into account  (Gosiewska and Biecek 2019; tsang2020does; mase2019explaining; Casalicchio, Molnar, and Bischl 2019; Hooker, Mentch, and Zhou 2021) A contradiction may thus appear between the semantic relationship of two notions, their actual correlation in the data used to train the model, and the explanation returned by the system. The second scenario is when XAI systems generate, rather than a single explanation, a set of explanations (e.g. of counterfactual examples (Mothilal, Sharma, and Tan 2020)) to provide richer insights to the user. In the counterfactual case, (Laugel et al. 2023) discuss the underlying notion of explanation diversity, which we connect here to the risk of getting competing explanations: diverse explanations generally aim to suggest alternative, hence sometimes contradictory, actions to the user.
Why is it a problem?
Whether coming from different explanations among a diverse set, or from diverse individual components within a given explanation, a contradiction between two pieces of information may be perceived as confusing by the user, potentially leading them to rejecting the explanation (Suffian et al. 2022). For instance (Bove et al. 2022) suggest competing explanations as one of the reasons for which users misunderstand the explanations returned by the SHAP approach. Furthermore, in a non-cooperative setting (Bordt et al. 2022) where the objectives of the user and the ML developer are not aligned, (Goethals, Martens, and Evgeniou 2023) discuss how this problem also opens up the risk of explanation manipulation through the selection of an explanation that is not in the best interest of the user.
Solutions.
The proposed solutions can be grouped in two categories: the first one focuses on better XAI systems, either by adapting them, e.g. to take into account correlations (Aas, Jullum, and Løland 2021), or enriching them, e.g. proposing to calculate, in addition to the usual feature importance vectors, the contributions of feature interactions (Gosiewska and Biecek 2019). A second type of enrichment consists in exploiting expert knowledge to contextualize feature contributions: informing the user, usually through the XUI, may allow to rationalize some confusion that can be caused by competing explanations (Bove et al. 2022).
Unstable explanations
We call explanations unstable when there is an inconsistency, i.e. a contradiction, between explanations within a supposedly stable scenario: for instance, when producing local explanations for similar instances with similar outcomes, one might expect that the explanations should be similar as well (Jacovi and Goldberg 2020). Such explanation inconsistencies have been widely observed, in particular for post-hoc XAI methods (Alvarez-Melis and Jaakkola 2018a; Laugel et al. 2019). Similarly to competing explanations unstable explanations are perceived originally as a technical failure of the explainer. Yet, as we describe below, it can also originate from the user, or from a combination of both the system and the user.
Why does it happen?
Unstable explanations are often considered as a technical failure of the explainer, viewed as a lack of robustness that needs to be fixed (Alvarez-Melis and Jaakkola 2018b; Slack et al. 2021; Mishra et al. 2021; Kindermans et al. 2019) However, other studies ascribe these inconsistencies to the model to be explained (Dombrowski et al. 2019; Ghorbani, Abid, and Zou 2019; Alvarez-Melis and Jaakkola 2018b): the local behavior of the latter may indeed vary abruptly, due to the complexity of the task being modeled, the complexity of the model itself or its lack of robustness. Faithful explanations then reflect these steep changes, leading to an apparent lack of stability. Moreover, this issue of instability may conflate with user perception of the similarity between instances and the explanations they expect as a result. This similarity, that depends on the user knowledge and possible biases, may differ from the similarity considered by the ML system. Pushing the expectation of explanation stability to its extreme, explanations generated for identical observations are anticipated to be identical. However, post-hoc model-agnostic methods, be they local or global (e.g.  (Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017; Wachter, Mittelstadt, and Russell 2018]; Altmann et al. 2010)), often rely on a stochastic data generation step (Zhou, Hooker, and Wang 2021; Visani et al. 2022) that may cause instability. This comes in addition to the Roshomon effect, i.e. that several equally good but potentially drastically different solutions can coexist and therefore be selected as explanations (Hancox-Li 2020).
Why is it a problem?
Considering that faithful explanations reflect the state of ML model, depending on its potential causes discussed above, a lack of stability can either be seen as an actual failure or as a desired characteristic: for instance more local explanations are expected to be less stable (Molnar 2020), with locality being a commonly expressed desideratum for explanations (Radensky et al. 2022). Thus, as for competing explanations, interpreting the lack of stability as a failure depends on the users needs, their knowledge of the explainer, and their perception of how similar the explanations should be: a user not being aware of the locality-stability tradeoff may see it as problematic but may not otherwise (Hancox-Li 2020). Still, instability may result in the user rejecting the explanation, or even the whole AI system, seeing it as a proof of unfairness by the model (Sharma, Henderson, and Ghosh 2020; Hickey, Di Stefano, and Vasileiou 2021), that may be abused by the organization providing the explanation (Goethals, Martens, and Evgeniou 2023).
Solutions.
Various approaches have been proposed to address the unstability issue, depending on its source. To fix the stochastic unstability of model-agnostic post-hoc explainers, most contributions focus on algorithmic modifications of the random data generation step they rely on, e.g. fully replacing it with a deterministic one (Zhou, Hooker, and Wang 2021; Zafar and Khan 2019), or through a reweighting strategy (Shankaranarayana and Runje 2019). In cases when unstability originates from a misalignment between the user perception and the ML system representation, several works propose strategies to constrain the ML model during its training phase (Alvarez-Melis and Jaakkola 2018b; Dombrowski et al. 2019). On a different note, rather than mitigating the problem, several works propose to measure the explanation stability, arguing that it describes a notion of uncertainty that can be helpful for the user to better understand and use them (Gosiewska and Biecek 2019; Shankaranarayana and Runje 2019; Slack et al. 2021; Visani et al. 2022).
Incompatible explanations: the ”Disagreement Problem” in XAI
A third case where several explanations can contradict each other occurs when several explainers are used in the same setting. We call such contradictions incompatible explanations, they correspond to the prevalent (Krishna et al. 2022) and well known Disagreement Problem (see e.g., (Sundararajan and Najmi 2020; Okeson et al. 2021; Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Neely et al. 2021)).
Why does it happen?
Assuming that the explanations are faithful and stable, the most high-level root cause of this issue comes from the fact that the task of providing explanations, in particular in the post-hoc setting, is essentially underdetermined (Bordt et al. 2022): as mentioned in the introduction, the concept of ”good” explanation is elusive and has been formalized in numerous different ways, taking into account different types of desiderata. For instance, because they rely on different assumptions, the explanations generated by LIME and SHAP are expected to differ, even if both are faithful (Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022). Similar discussions apply to global feature attribution methods (Okeson et al. 2021) or counterfactual explanations (Goethals, Martens, and Evgeniou 2023). Going further, some differences between explanations can also be attributed to discrepancies in the implementations of the supposedly same mathematical explanation objective (e.g. (Sundararajan and Najmi 2020) for Shapley value-based explanations). Finally, a common source of incompatibility could be attributed for the explanations generated using the same method, but different parameters. Indeed, XAI methods generally rely on hyperparameters that may not always be understood (if known at all) by the user, albeit heavily impacting the obtained explanations (Garreau and Luxburg 2020). As a result, explanations can differ on multiple bases, ranging from the top features being different to differences in order of importance or direction (Krishna et al. 2022).
Why (and when) is it a problem?
This disagreement between explanations is generally viewed as a problem (Sundararajan and Najmi 2020; Garreau and Luxburg 2020; Poyiadzi et al. 2021; Swamy et al. 2022; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Roy et al. 2022; Goethals, Martens, and Evgeniou 2023). Several user studies have noted it to be a source of confusion for users (Okeson et al. 2021; Krishna et al. 2022), resulting in their lowering trust in the system and therefore leading to a reject of the system as a whole or a questionable selection of the proposed explanations (e.g. based on method popularity (Kaur et al. 2020)). As for competing explanations, in a non-cooperative setting, this incompatibility between explanations may also be seen as a degree of freedom that a malicious AI practitioner may leverage to rationalize unfair decisions by choosing the explanation the most aligned to their objectives  (Slack et al. 2020; Aïvodji et al. 2019; Goethals, Martens, and Evgeniou 2023).
On the other hand, similarly to the unstable and competing cases, incompatible explanations, if faithful, can be viewed as an opportunity that may be leveraged for a better interaction with the system in a collaborative context. Indeed, explanation disagreement can be seen as a source of diversity, as noted by (Laugel et al. 2023; Goethals, Martens, and Evgeniou 2023), which is viewed positively and helps understanding, as empirically shown by (Bove et al. 2022) when combining counterfactual with global feature attributions. In other cases, disagreements between feature importance explanations are seen as an evidence of a lack of robustness on the side of the classifier, therefore used as a starting point for model auditing (Okeson et al. 2021) to improve its performance.
Solutions.
Intuitively, circumventing the issue of incompatibility may be simply done by hiding disagreements (Roy et al. 2022). On the contrary, other works propose to emphasize them through interfaces (Bove et al. 2023), sometimes arguing, like for stability, that the level of disagreement may be used as a measure of uncertainty to validate parts of the explanations (Krishna et al. 2022). Using the same idea, other works propose to aggregate various explanation methods (Bhatt and Moura 2021; Pirie et al. 2023) to provide more robust explanations, or even propose to train new models that minimize this disagreement (Schwarzschild et al. 2023).
User-specific failures
This section discusses explanation failures that can occur when the users misinterpret the explanations provided by a ML system with no technical failures: regardless of their quality, these explanations have been shown to sometimes fail in their explanatory objective (Cheng et al. 2019; Wang et al. 2019). We propose here a typology of failures that originates from inconsistent users’ inferences, distinguishing between three categories graphically represented in Fig. 2 and discussed in turn below: mismatch failures when there is a contradiction between the ML explanations and the users’ expectations in term of format; counterintuitive failures when this contradiction concerns the explanation content; and biased inferences failures when cognitive biases inherent to each user interfere with the explanations. We believe that these user-specific failures should be known so XAI designers can understand the users’ mental model processes and support their interpretation of the provided explanations.
Figure 2: Types of user-specific explanation failures in ML explanations. The ML model is supposed to give an accurate prediction and the explainer to generate faithful explanations. The cross indicates the step at which the failure takes place.
Mismatch
First, we propose to define mismatched explanations when the format of the extracted information does not meet the users’ expectations towards the ML system (i.e., ”This is not what I want” in Fig. 2).
Why does it happen?
Depending on the context of the interaction and the nature of the decision model’s outputs, users have been shown to have different needs and questions regarding the ML system and the user questions may also vary depending on the model’s output  (Liao, Gruen, and Miller 2020; van der Waa et al. 2021). In parallel, there is a huge diversity in the forms explanations can take (see e.g. (Dwivedi et al. 2023)), but explanation techniques do not necessarily take into account the context in which the explanations is sought by a user (Matarese, Rea, and Sciutti 2021). As a consequence, mismatches between the user questions and the generated explanations can occur, on the explanation type (e.g., feature importance or rules), locality (e.g., local or global), goal (e.g., factual or causal) and on the information complexity (e.g., expressed for ML practitioners or lay users), to name a few.
Such mismatches reflect the lack of user-centricity in the conception of the XAI system. Actually, most XAI approaches are designed without evaluating whether the explanations satisfy the needs of real users (Miller 2019). Instead, other criteria are used to evaluate the relevance of an XAI approach, such as the visual aspect of the explanations, its popularity or the ease of implementation (Mohseni, Zarei, and Ragan 2018; Barredo Arrieta et al. 2020). Such evaluations can bring to light that they fail to meet these needs: for example, during the co-design workshop for an AI-based diagnosis tool (Wang et al. 2019), many interviewed doctors reported that they would prefer alternative hypotheses (e.g., counterfactual examples) rather than factual explanations (e.g., local feature importance technique) that were initially implemented in the offered system.
Why is it a problem?
Mismatched explanations can lead to dissatisfaction and a reject of the overall ML system: It has been demonstrated that differences between initial expectations and actual experiences can affect both the user satisfaction and acceptance of a system (Bhattacherjee 2001). Along the same lines, the complexity of some explanation types can hinder users to understand the explanations (e.g., if they lack skills in AI), and thus the adoption of the whole ML system (see e.g. (Kaur et al. 2020) in the case of SHAP and GAM). It has also been demonstrated that users assign artificial agents human-like traits, and hence expect these agents to provide explanations using the same conceptual framework they are used to (De Graaf and Malle 2017). Yet, most XAI solutions proposed so far have not reached such a complete human-centred approach (Miller 2019; Matarese, Rea, and Sciutti 2021).
Solutions.
Mitigating the mismatched explanations somehow obviously relies on applying user-centered approaches when designing XAI systems: two examples include adjusting the types of explanations according to users’ expectations to improve their satisfaction and acceptance (Riveiro and Thill 2022) or integrating a lexical alignment step to improve the understanding of explanations provided by a conversational agent (Srivastava, Theune, and Catala 2023). Another type of solution aims at integrating social and cognitive science’s theories in a theoretical framework for XAI system, to provide personalized and contextualized explanations (Matarese, Rea, and Sciutti 2021).
Counterintuitive explanations
We then propose to identify counterintuitive explanations when the explanation provided by the ML system is in contradiction with the prior knowledge of expert users (i.e., AI practitioners and domain experts). As compared to the mismatch failure for which explanations can be contradictory with users expectations in terms of format, counterintuitive failures can be observed when the content of the explanation itself is contradictory with users’ knowledge.
Why does it happen?
Prior knowledge may take various definitions (Dochy and Alexander 1995), we view it here as ”stored knowledge about the world that have been acquired by an individual” (Brod, Werkle-Bergner, and Shing 2013), including domain expertise and past experience. Now in an XAI context, some users can find faithful explanations contradicting with such prior knowledge, leading them to perceive these explanations as different from what they anticipated (”This is not what I expected” in Fig. 2). This is all the more likely to happen as explanations are especially requested when the model’s output is perceived by users as abnormal or absurd (Riveiro and Thill 2022). In the same vein, past experience has been shown to lead to disagreement with the explanations (Sohn et al. 2019; Suffian et al. 2022).
Why is it a problem?
Counterintuitive explanations do not necessarily represent a failure: explanations that do not match user expectations can indeed be used to fix potential issues within the ML system (Kaur et al. 2020) or for knowledge discovery (Jiménez-Luna, Grisoni, and Schneider 2020). In other situations however, perceiving the provided explanations as counterintuitive can lead users to question the reliability of the prediction even when it is accurate (e.g., (Collaris, Vink, and van Wijk 2018) in the context of insurance fraud and (Palaniyappan Velumani et al. 2022)in the context of environmental studies), or of the ML system as a whole (e.g. (Nourani et al. 2021)). They can even perceive the explanations as being absurd if the latter do not correspond to their past experiences (Sohn et al. 2019). This aligns with findings from works in social sciences, which have shown that people tend to ignore an information that is inconsistent with their beliefs from past experiences (Thagard 1989; Nickerson 1998).
Solutions.
Mitigating these failures requires to better align ML explanations and users’ prior knowledge. Some works in XAI propose to integrate the users’ knowledge directly in the explanation generation (e.g., (Jeyasothy et al. 2022)). Conversely, a different approach proposes to change the predictive model itself to facilitate the generation explanations that are aligned with users’ prior knowledge (Rieger et al. 2020). In a different perspective, another work proposes to co-design explanation interfaces together with expert in co-design workshops so as to integrate both their needs and knowledge (Wang et al. 2019).
Biased inferences
Finally, we propose to identify explanation failures relying on biased inferences when users make inaccurate interpretations of the explanations, due to cognitive biases (see Fig. 2). As compared to counterintuitive failures that occur for expert users, for biased inferences we consider mainly lay users that do not have expert knowledge nor past experiences. Yet, such failures can occur for any type of users as cognitive biases are inherent to all humans.
Why does it happen?
Similarly to prior beliefs, biases can influence how users respond to different styles of explanations (Liao, Gruen, and Miller 2020; Miller 2019), leading eventually to their inaccurate interpretation (Bertrand et al. 2022; Nourani et al. 2021; Wang et al. 2019). We discuss below some common cognitive biases with their consequences on the interpretation of explanations.
Why is it a problem?
Several cognitive biases have been shown to trigger inaccurate interpretations of explanations (Bertrand et al. 2022). First, some biases can trigger over reliance in ML explanations. It has been demonstrated that having an explanation, regardless of its quality, increases trust (Eiband et al. 2019; Lai and Tan 2019). In other examples, it is shown that longer explanations are found to be more plausible than shorter ones (Fürnkranz, Kliegr, and Paulheim 2020), and that users may believe they understand better than what they actually do (Rozenblit and Keil 2002; Mueller et al. 2019; Chromik et al. 2021). On the other hand, other biases can trigger under reliance. For instance, the ”negativity bias” can cause lay users, in particular, to pay more attention or overweight negative information over positive one of the same strength (Kliegr, Bahník, and Fürnkranz 2021). It may lead users to pay more attention to negative outcomes of the ML system, thus eroding their trust (Pratto and John 1991). It has been demonstrated that showing the weaknesses of the system (e.g., competing explanations) or negative outcomes (e.g., such as a malignant diagnosis) early on can have a major influence on trust (Nourani et al. 2021). Finally, some other biases can trigger users to misapply the explanations: e.g., the ”insensitivity to sample size” bias may lead lay users to ignore the statistical significance of a statement (Fürnkranz, Kliegr, and Paulheim 2020); e.g., the ”availability” may lead lay users to believe that examples and events that easily come to mind are more representative than is actually the case (Wang et al. 2019); e.g., the ”primacy effect” bias may lead them to form an opinion based solely on the first piece of information received (Nourani et al. 2021).
Solutions.
Before mitigating these biased inferences, identifying them and measuring their effects on the users’ interpretation is a challenging task that can e.g. rely on comparing users’ objective and self-reported understanding (Cheng et al. 2019; Bove et al. 2022). Most approaches then rely on the design of appropriate XUI (Wang et al. 2019), for instance controlling what types of predictions users first see when interacting with the system, to mitigate the negative bias (Nourani et al. 2021).
Discussion and research agenda
The typology of explanation failures presented in the previous section allows to understand why failures happen, how to mitigate them, and how to distinguish them from one another. In this section, we leverage this typology to discuss some key issues, and identify some promising research directions.
Towards a holistic XAI approach
Observation: Some failures result from the interaction between components of the explanation process.
Many of the explanation failures discussed in the previous sections can be diagnosed as stemming from one of these components (model, explainer and user): they can for instance be due to a technical problem with the system. However, our analysis in the previous section also underlines that some errors actually arise from the interplay between the different components considered, rendering them incompatible: we studied the interaction between the explainer and the user in the third section about user-specific failures and discussed in the fourth section some cases of problematic interaction between the model and the explainer. For instance, issues like unstability may highlight a mismatch between the decision model’s behavior (volatile, local), the implicit assumptions made during explanation design (stable boundaries) and the similarity between instances perceived by the user.
Consequence: subpar technical solutions.
In the case of these interaction between components, the failure is not caused by a deficiency in either component, but rather from their misalignment. Consequently, some technical solutions to address these failures suggest adapting or replacing one of the components afterwards, sometimes taking a paradoxical turn: instead of questioning the explainer when observing a failure, it is sometime envisaged instead to train a new AI decision model that would be more adapted to this explainer. As an illustrative example, it has been proposed to build models with more stable behavior to mitigate instability issues of gradient-based explainer (Alvarez-Melis and Jaakkola 2018b), or models constrained to minimize the disagreement between LIME and SHAP (Schwarzschild et al. 2023). This may seem surprising, as explainers are usually leveraged to generate insights about the model, not the other way around.
Going forward: towards a holistic design of the explanation process.
This underscores a significant challenge: the interplay between the elements of the system should be taken into account from the system’s inception. The three components of the system should thus be regarded as interconnected, rather than being designed independently. This does not mean abandoning post-hoc methods but rather anticipating their integration in the overall explanation process.
Multiple calls for a user-centric approach of XAI have been made, pushing for taking the user needs into account from the inception of the system and guiding the design of the XAI method (Wang et al. 2019). Nevertheless, more efforts should be pursued in studies on the interaction between AI models and explainers. One possibility is to draw inspiration from research in Integrative Design for software systems (Tumer and Smidts 2010) which proposes holistic design strategies for software systems. Design and monitoring of AI systems should thus be conceived in a holistic way, with any choice or change in the decision model prompting a reassessment of its compatibility with the explainer, and vice versa; and similarly for changes in user requirements.
Towards less one-size-fits-all and more transparent XAI solutions
Observation: some failures happen because explainers are black-boxes, and both ML practitioners and users do not know their limitations.
The previous argument about interaction failures and the need to adopt a holistic approach of AI systems design also raises the question of the relevance of one-size-fits-all XAI solutions. The most well-known XAI methods (e.g. SHAP and LIME) are often conceived under assumptions of data-, model-, and user-agnosticity (i.e. absence of user focus), generally with the aim of providing more flexibility in their use. Yet, some of the failures discussed in the third section about system-based failures and the previous argument about interaction failures suggest that these methods may not, in fact, be one-size-fits-all solutions. This adds up to the existing research already questioning the often claimed user-agnosticity, suggesting that some types of explanations may in fact not be suited to all user profiles (Wang et al. 2019)), or to all decision models (Alvarez-Melis and Jaakkola 2018b). This restates that explainers face various limitations and various assumptions, that should be known and understood for proper use.
Consequence: ML practitioners and users misuse XAI systems.
Unfortunately, these limitations and assumptions of XAI systems are rarely known to users and even ML developers, as analyzed in the typology we propose. While this may seem intuitive for user-specific failures (e.g. overtrust issues in biased inferences), it is also true for system-based failures: for instance, some competing issues are caused by the user not knowing how the explainer handles correlations and interactions between features; some instability issues by the user not understanding the tradeoff between stability and locality, nor what level of locality they would like; some incompatibility issues by the user not understanding the differences actually captured by two explainers. Consequently, users may reject or misuse the explanation, not because it does not meet their needs, but because they do not understand how to use it, as shown empirically by  (Kaur et al. 2020) for SHAP.
Going forward: more ”transparency” in XAI methods.
Technical design choices of XAI systems and the assumptions they rely on heavily impact how the explanation should be interpreted. Overall, this further confirms the need for more effort in communicating on the core capabilities of XAI systems, e.g. through design principles such as ML transparency (Bove et al. 2022). This aligns with prior calls to further encourage more interdisciplinary works for designing more transparent XAI systems, instead of building black-box explainers. However, research in this remains timid and more efforts should be made. One direction to pursue is the possibility of describing explanation methods in a factual way, similarly to Model Cards (Mitchell et al. 2019).
Towards more XUI
Observation: some failures happen because the users are limited in their interaction with the explanations.
From a cognitive point of view, the explanation process is argued to be interactive (Miller 2019) and yet, at best, it is merely sequential in XAI. Most approaches are limited to the generation of factual information about the model’s behaviour or the predicted outcome (e.g., feature importance scores, counterfactual examples, etc.) and the users are not always able to interact with it.
Consequence: external information and processes interfere with the explanations.
This causes the explanations to be potentially perceived by users as being incomplete or incorrect (e.g., lack of knowledge or lack of transparency on the XAI method). As discussed in the previous section, users may thus draw on external information and cognitive processes to interpret the explanations, leading to potentially one or more XAI failures. For instance, users may expect that there can be intuitive changes in counterfactual examples because they have experienced the same logical path in real life (Suffian et al. 2022).
Going forward: more user interfaces for explanations.
We believe that XUIs allow to better organize, complete and display ML explanations according to the user needs (Chromik and Butz 2021). Moreover, accounting for the users is key to design such interfaces, which thus forces AI practitioners to adopt a user-centered approach when conceiving ML systems (see Section Observation: Some failures result from the interaction between components of the explanation process.). Previous studies have demonstrated the usefulness of visual interfaces to present ML explanations (e.g., (Szymanski, Millecamp, and Verbert 2021; Ooge, Kato, and Verbert 2022)) but there are other interaction modalities, in particular the conversational mode. As users have progressively become more familiar with conversational AIs (e.g. ChatGPT), the provided explanations may thus be presented in conversational interfaces and become a real dialog, i.e. questions from the users and corresponding answers from the machine (Miller 2019). This could open new perspectives for design principles such as introducing a narrative logic that allows temporizing the current information. For instance (Nourani et al. 2021) suggest to control primary information to mitigate negativity bias. We believe that studying such a modality for the display of explanations would allow to better understand users’ processes for analyzing and understanding ML explanations.
Conclusion
In this work, we have proposed a typology of XAI failures allowing to understand why failures happen, how to mitigate them, and how to distinguish them from one another. We believe it can help AI developers and designers better understand XAI systems and their limitations. Leveraging this typology, we have identified promising research avenues for XAI. Besides these directions, future works include investigating the potential combinations of multiple failures and their consequences on user understanding. Furthermore, a better understanding of each failure could be leveraged to formally propose a diagnostic framework to help identifying their origins in the explanation process.