### v2

# Why Do Explanations Fail? A Typology and Discussion on Failures in XAI

Abstract
As Machine Learning models achieve unprecedented levels of performance, the XAI domain aims at making these models understandable by presenting end-users with intelligible explanations. Yet, some existing XAI approaches fail to meet expectations: several issues have been reported in the literature, generally pointing out either technical limitations or misinterpretations by users. In this paper, we argue that the resulting harms arise from a complex overlap of multiple failures in XAI, which existing ad-hoc studies fail to capture. This work therefore advocates for a holistic perspective, presenting a systematic investigation of limitations of current XAI methods and their impact on the interpretation of explanations. By distinguishing between system-specific and user-specific failures, we propose a typological framework that helps revealing the nuanced complexities of explanation failures. Leveraging this typology, we discuss some research directions to help practitioners better understand the limitations of XAI systems and enhance the quality of ML explanations.
1 Introduction
The field of eXplainable Artificial Intelligence (XAI) aims at addressing the challenge of providing users with explanations regarding decisions of Machine Learning (ML) models, bridging the gap between the inner workings of complex algorithms and human understanding. It constitutes a multidisciplinary domain, drawing upon not only computer science but also cognitive sciences, philosophy and human-computer interaction (Miller 2019; Byrne 2023; Liao, Gruen, and Miller 2020; Molnar 2020; Zednik 2021).
Central to the study of XAI is the elusive concept of a ”good” explanation. Formal definitions, see e.g. (Amgoud and Ben-Naim 2022), do not capture the human component and prove to be a daunting task, leading researchers to gradually develop ad-hoc desiderata and investigate specific challenges that arise. The domain multidisciplinarity often results in fragmented investigations, with different research communities focusing on disparate aspects of the problem. On one side, some limitations of explainers such as their lack of robustness (Alvarez-Melis and Jaakkola 2018a) or faithfulness (Jacovi and Goldberg 2020; Lyu, Apidianaki, and Callison-Burch 2024) have been mostly investigated by the ML community. In parallel, other works in different fields (e.g. social sciences) have pointed out issues such as the difficulty for explanations to meet user needs (Matarese, Rea, and Sciutti 2021), prior beliefs (Riveiro and Thill 2022) or general reasoning (Bertrand et al. 2022). However, this fragmented approach poses a significant challenge, as the limitations that XAI systems face are not mutually exclusive; we argue that they may overlap and conflate with each other. Consequently, AI practitioners may find themselves unable to comprehend the origin of a failure, and thus to properly mitigate the resulting harms.
To circumvent this issue, we argue that it is necessary to go beyond existing ad-hoc and domain-specific discussions on XAI issues and adopt a holistic approach to XAI failures. Leveraging existing works on XAI approaches, interfaces and evaluations, the main contribution of this paper is the first typology that encompasses insights from both the XAI-ML and the XAI-HCI (Human-Computer Interaction) communities, thereby directly accounting for the complex multidisciplinarity of the XAI field. We provide a systematic and hierarchically organized overview of XAI failures, distinguishing between system-specific and user-specific ones. Contrary to a systematic literature review, our goal is not to cover all existing works, but rather discuss, for each failure, their origin, characteristics, and some potential mitigation solutions from the literature. We believe that this typology can help AI practitioners gain a deeper understanding of their connections and origins. Leveraging this typology, we then bring together the system-centric and user-centric perspectives to discuss research avenues to enhance the quality of explanations provided by XAI systems. By fostering a more nuanced understanding of the limitations inherent to XAI, we hope to pave the way for more effective and transparent automated decision-making processes.
This paper is organized as follows: after describing in Section 2 the context and the relation to existing works, we discuss in Section 3 our research objectives, as well as the methodology followed to build the proposed typology of XAI failures. The latter is then presented in two sections corresponding to its root split: Section 4 discusses the system-specific failures and Section 5 the user-specific ones. Finally, we discuss in Section 6 insights provided by the typology, answering our research questions.
2 Context and Related Works
This background section successively discusses the setting we consider and related works.
2.1 Explanation Process Components
The explanation process is generally seen as being composed of two components of different nature: the Machine Learning system on one hand and the user on the other hand.
The Machine Learning system (ML system) is in turn composed of two components, that may be entangled and difficult to distinguish: the ML model, that provides decisions, and the explanation method, that generates rationale for these predictions. We consider a supervised learning context, where given some input information, a ML model returns an associated decision. The data used as input can be either structured (e.g. tabular) or unstructured (images, text, etc.). The nature of the model can vary over a wide range, from simple (e.g. linear) models to deep neural networks or large language models. The model performance can be evaluated using various metrics of performance, e.g. accuracy or computational complexity to name two examples. In addition to the prediction itself, the system provides rationale for it, through an explanation generator. It can either be the predictive model itself (in the case a transparent model is used) or a separate system, composed of one (or several) explainer(s), built on top of the predictive model. There exists a huge diversity of methods to generate various kinds of information that act as explanations (see e.g. Dwivedi et al. (2023) for a recent survey), either for one prediction (local explanation) or for the whole model behavior (global explanation). Additionally, several works propose to design eXplanation User Interfaces (XUIs), see e.g.  Chromik and Butz (2021), to display the generated explanation to the end-users in an intelligible and useful manner.
The counterpart within this two-part explanation process is the user who receives the explanation and interacts with the ML system to accomplish their task. The explanation should allow them to understand the decisions made by the model. It must be underlined that users can have various objectives, e.g. depending on their expertise levels and prior knowledge (Liao, Gruen, and Miller 2020), to name two examples, that can lead to different needs in terms of interpretability, see e.g. (Mohseni, Zarei, and Ragan 2018).
Notion of Explanation Failure
As a consequence of this explanation process structure, it can be argued that a successful explanation depends on three elements: the ability (i) of the ML model to make an accurate prediction, (ii) of the explainer to provide a faithful explanation that addresses user needs, and (iii) of the user to properly understand and use it. When at least one of these elements fails, we say that there is an explanation failure that needs to be investigated.
2.2 Related Works
There is an abundant literature focusing on XAI limitations, that we detail in the next sections. However, the vast majority of these works address such failures by adopting a technical point of view, as opposed to a human-in-the-loop approach, see e.g. Molnar et al. (2020); Barredo Arrieta et al. (2020); Srivastava et al. (2022); Saeed and Omlin (2023); Bodria et al. (2023) for some overviews. In comparison, very few works suggest that limitations stemming from the user side should also be investigated, such as mismatches between explanations and user needs (Matarese, Rea, and Sciutti 2021) or cognitive biases (Bertrand et al. 2022).
In addition, most existing works consist in ad hoc studies on specific problems, viewing them as independent from one another. This is further exacerbated by the fact that technical failures and issues on the user side of the explanation process are generally studied in different domains (Computer Science for the first, Human Computer Interaction and Cognitive Sciences on the other). Yet, given the interactive nature of the explanation process (Hilton 1990), it is likely that addressing issues in a more global manner is needed: the process of an explanation is sequential, a ”conversation” between the ML system first providing predictions and explanations and then the user interpreting and possibly interacting with them (Miller, Howe, and Sonenberg 2017; Miller 2019). It can therefore be expected that some failures may interact, conflate, or even amplify one another, raising the need for a holistic perspective on XAI issues. While some contributions in this direction have been proposed, they remain generally focused on domain-specific contexts (Vellido 2020; Antoniadi et al. 2021).
3 Research Questions and Methodology
This section describes the methodology we implement to build the proposed typology on XAI failures, after discussing the research questions we identify.
3.1 Research Questions
As stated in the previous sections, the main research question we address can be formulated as follows:
RQ1: What are the different failures that may arise during the explanation process?
A related valuable information concerns the risks the failures can lead to. In this paper, we do not consider the issue of malicious uses of explanations or deliberate intent to fool explanation systems, and the domain of deceptive XAI (Dimanov et al. 2020; Lakkaraju and Bastani 2020; Slack et al. 2021a; Schneider, Meske, and Vlachos 2023), to be a an explanation failure in itself. However, we include a discussion on the explanation dysfunctions that may be exploited by potential malicious actors.
The identification, and structuration, of failures can help on the way to propose detection and mitigation strategy, which leads to the following second research question.
RQ2: How can these failures be avoided?
To answer this question, it is crucial to understand how failures happen and possibly the reasons why they can occur. The typology we propose thus includes discussions regarding these topics.
In order to answer these questions, we apply a paper-guided approach, following the methodology described in the next subsection, basing the proposed typology on the analysis of related publications. However, contrary to a systematic literature review, our goal is not to cover all existing works and to provide an exhaustive survey, but to propose a categorization of failures identified in the literature.
Meta-characteristic Failure name
Discuss the failure or its consequences (Why it happens? and Why is it a problem?)
Discuss solutions
System-specific Misleading
Laugel et al. (2019); Ye and Durrett (2022); Papenmeier, Englebienne, and Seifert (2019)
Jacovi and Goldberg (2020); Laugel et al. (2018b); Han et al. (2023)
Kaur et al. (2020); Agarwal, Tanneru, and Lakkaraju (2024); Colin et al. (2022)
Jacovi and Goldberg (2020); Laugel et al. (2018b)
Han et al. (2023); Agarwal, Tanneru, and Lakkaraju (2024)
Li et al. (2023)
Competing
Gosiewska and Biecek (2019); Tsang, Rambhatla, and Liu (2020)
Casalicchio, Molnar, and Bischl (2019); Hooker, Mentch, and Zhou (2021)
Suffian et al. (2022); Bove et al. (2022); Laugel et al. (2023); Zhou et al. (2021)
Goethals, Martens, and Evgeniou (2023); Zhou and Joachims (2023)
Mase, Owen, and Seiler (2019); Mothilal, Sharma, and Tan (2020)
Aas, Jullum, and Løland (2021); Bove et al. (2022)
Gosiewska and Biecek (2019); Salih et al. (2024)
Jiang et al. (2025)
Unstable
Jacovi and Goldberg (2020); Alvarez-Melis and Jaakkola (2018a); Hancox-Li (2020)
Slack et al. (2020); Mishra et al. (2021); Kindermans et al. (2019)
Dombrowski et al. (2019); Ghorbani, Abid, and Zou (2019); Zhou, Hooker, and Wang (2021)
Sharma, Henderson, and Ghosh (2020); Molnar (2020); Radensky et al. (2022)
Visani et al. (2022); Hickey, Di Stefano, and Vasileiou (2021); Laugel et al. (2019)
Zhou and Joachims (2023); Goethals, Martens, and Evgeniou (2023)
Zhou, Hooker, and Wang (2021); Zafar and Khan (2019)
Alvarez-Melis and Jaakkola (2018b); Slack et al. (2021b)
Dombrowski et al. (2019); Visani et al. (2022)
Gosiewska and Biecek (2019); Yeh et al. (2019)
Shankaranarayana and Runje (2019)
Incompatible
Krishna et al. (2022); Okeson et al. (2021); Bansal, Agarwal, and Nguyen (2020)
Bordt et al. (2022); Neely et al. (2021); Goethals, Martens, and Evgeniou (2023); Kaur et al. (2020)
Swamy et al. (2022); Roy et al. (2022); Slack et al. (2020); Reingold, Shen, and Talati (2024)
Aïvodji et al. (2019); Laugel et al. (2023); Bove et al. (2022); Garreau and Luxburg (2020)
Sundararajan and Najmi (2020); Han, Srinivas, and Lakkaraju (2022); Poyiadzi et al. (2021)
Roy et al. (2022); Bove et al. (2023); Krishna et al. (2022)
Pirie et al. (2023); Schwarzschild et al. (2023)
Bhatt and Moura (2021); Decker et al. (2024)
User-specific Mismatch
Liao, Gruen, and Miller (2020); van der Waa et al. (2021); Dwivedi et al. (2023)
Doshi-Velez and Kim (2017); Miller (2019); Mohseni, Zarei, and Ragan (2018)
Bhattacherjee (2001); Kaur et al. (2020); De Graaf and Malle (2017); Keane et al. (2021)
Barredo Arrieta et al. (2020); Wang et al. (2019); Matarese, Rea, and Sciutti (2021)
Srivastava, Theune, and Catala (2023); Zarlenga et al. (2024)
Matarese, Rea, and Sciutti (2021); Byrne (2023)
Riveiro and Thill (2022); Pazzani et al. (2022)
Counter-intuitive
Riveiro and Thill (2022); Sohn et al. (2019); Kaur et al. (2020); Nourani et al. (2021)
Jiménez-Luna, Grisoni, and Schneider (2020); Collaris, Vink, and van Wijk (2018)
Thagard (1989); Ebermann, Selisky, and Weibelzahl (2023); Dochy and Alexander (1995)
Brod, Werkle-Bergner, and Shing (2013); Nourani et al. (2021); Suffian et al. (2022)
Cabitza et al. (2024); Palaniyappan Velumani et al. (2022); Nickerson (1998)
Jeyasothy et al. (2022); Rieger et al. (2020)
Wang et al. (2019); Lim et al. (2025)
Conati et al. (2021); Ross, Hughes, and Doshi-Velez (2017)
Ebermann, Selisky, and Weibelzahl (2023); Koh et al. (2020)
Biased Inferences
Hoff and Bashir (2015); Liao, Gruen, and Miller (2020); Miller (2019)
Eiband et al. (2019); Lai and Tan (2019); Rozenblit and Keil (2002)
Chromik et al. (2021); Kliegr, Bahník, and Fürnkranz (2021); Pratto and John (1991)
Nourani et al. (2021); Bertrand et al. (2022); Mueller et al. (2019)
Wang et al. (2019); Fürnkranz, Kliegr, and Paulheim (2020)
Cheng et al. (2019); Wang et al. (2019); Bove et al. (2022)
Nourani et al. (2021); He, Kuiper, and Gadiraju (2023)
He, Aishwarya, and Gadiraju (2025)
Table 1: List of the references chosen to illustrate the typology, categorized by failure and type of contribution. Some references appear in several cells of the table.
3.2 Methodology
We build the proposed typology using the guidelines developed by Nickerson, Varshney, and Muntermann (2013) for Information Systems taxonomies, made of 5 steps: (1) Define a meta-characteristic, (2) Specify ending conditions, (3) Identify a subset of objects, (4) Identify common characteristics and group objects, (5) Group characteristics into dimensions to refine typology. Steps 3 to 5 are repeated iteratively until the ending conditions specified in step 2 are met. Steps 3 and 4 can be done in this order, called empirical-to-conceptual process, or in the reverse one, called conceptual-to-empirical, where 4 is rephrased as ”Conceptualize characteristics and dimensions of objects” and 3 as ”Examine objects for these characteristics and dimensions”. Following this principle, we alternate inductive categories extraction from papers and deductive categorization of papers.
Meta-characteristic
The meta-characteristic aims at providing a basis for identifying the other dimensions that the typology will rely on. All following characteristics are then intended to be logical consequences of the meta-characteristic, itself deriving from the research questions and the typology’s intended use. As the typology we propose to build aims primarily to cover XAI failures by adopting a holistic perspective covering both the ML system and the user, we use a binary meta-characteristic distinguishing system-specific failures, grouping issues associated to technical limitations of the ML system, from user-specific ones, which encompass issues caused by the inferences users make about the provided explanations.
Ending conditions
We use both objective and subjective criteria proposed by Nickerson, Varshney, and Muntermann (2013): the process stops when no new dimension or characteristic has been added in the last iteration. In addition, it stops when the typology is assessed to be concise (at most 10 types of failures), robust (at least 5 papers in each category), comprehensive and explanatory (the categories are easily distinguishable based on the characteristics).
Data collection
A crucial step is the collection of works relevant to the topic of XAI failures. This task has been performed in an iterative manner, enriching the set of collected papers through enriched list of search keywords. Included in the screening scope are the proceedings of the main venues from the fields of AI (ICML, NeurIPS, IJCAI, etc.), HCI (CHI, IUI, etc.), and explainability specialized conferences (FAccT, AIES, XAI conference, etc.). Were also considered papers available on ArXiv to scan for potentially unpublished but meaningful contributions.
The initial list of search keywords included terms as explainability/explanations/explaining, interpretability/interpreting, transparency, with and without associations with notions such as failures, problems, risks, pitfalls, inconsistencies, etc. Iteratively, after identifying new categories in the taxonomy, it was enriched through category-specific keywords such as stable/unstable/robust explanations, etc.
In all iterations, the inclusion criterion of the retrieved papers in the collection imposes that their contributions : (i) identify and discuss pitfalls of existing explainability methods, either from a theoretical or an empirical perspective; (ii) or propose new explanation methods to mitigate specific issues, with quantitative assessments of these results.
After summarizing the contributions of each paper and documenting the rationale behind their relevance for the typology, the authors discussed together their inclusion.
3.3 Overview of the Result
The methodology described in the previous section lead to select a total number of 108 papers to build the typology. After the typology was built, we check that each type, and in particular each leaf type, is associated with at least 5 papers, so as to ensure it is representative and significant.
The selection of considered characteristics and dimensions is derived from the considered meta-characteristic and driven by the considered research questions, related to the aim of avoiding these failures. For system-specific failures, discussed in Sect. 4, a temporal dimension related to the explanation process is taken into account, to define subtypes depending on the system development step at which dysfunctional behaviors may occur: the ML model itself, the explanation generator or the generated explanation that may contain conflicting pieces of information. The latter is further decomposed, at a third level of the typology, depending on the source of the conflict. For user-specific failures, discussed in Sect. 5, the structuring dimensions we propose distinguish whether the explanation is rejected or accepted by the user and additionally examine the rejection cause, depending on whether it related to the explanation form or content. In case of acceptation, a failure can occur in cases where the explanation is actually misunderstood or misused.
In addition, in order to answer more accurately the considered research questions, we propose to enrich each explanation failure type with a discussion along three axes: (i) why does the failure happen, (ii) why is it a problem and (iii) what kind of solutions, if any, have been proposed on the literature, either to measure or inform about the issue, mitigate its negative consequences or even solve it. Regarding (ii), it can indeed be observed that, depending on the context, a phenomenon can be seen as an issue or not. This can e.g. be related to the fact that even explanation manipulation can be seen as desirable in specific cases: Slack et al. (2020) argue that it can be used as as a method to preserve intellectual property about the classifier, avoiding to disclose its underlying principle. An overview of the typology, with the references considered to support it, is shown in Table 1.
4 Proposed Typology: System-specific
XAI Failures
Figure 1: Types of system-specific explanation failures: misleading explanations ascribed to the ML system and contradicting ones ascribed to the explainer. The cross and the $\neq$ symbol indicate the step at which the failure takes place.
This section discusses explanation failures that can be ascribed to the Machine Learning system, depending on its development step at which they can occur. A graphical representation of the 4 proposed subtypes, organized in two categories, is provided in Fig. 1 and commented below.
4.1 Overview
The typology decomposes system-specific explanation failures into two categories: (1) misleading explanations when either the ML model provides an inaccurate prediction or when the explainer is not faithful; and (2) contradictions when conflicting information are provided by one or several explainers. The effect of the former on the users can be characterized by the summarizing question ”I understand the explanations, but should I?”.
The latter can be further decomposed into 3 categories depending on the source of the conflict: inconsistencies can occur because of contradiction between (a) different pieces of information of the same explanation, leading to explanations we propose to name competing, (b) different explanations generated by the same explainer, named unstable explanations, or (c) different explanations generated by different explainers, named incompatible explanations. This case may occur when the global explainer is defined as a set of explainers. In other words, as the diagram in Fig. 1 illustrates, the plurality that leads to the contradiction can be due to the output, the input or the explainer itself. In all three cases, users may not understand this conflicting information, and wonder ”Why is it different?”
4.2 Misleading Explanations
We call an explanation misleading when the failure results from the ML system being dysfunctional, i.e. when it fails to meet the very purpose it was designed for. We distinguish two situations, depending on whether the dysfunction comes from the prediction or the explanation. There is a risk that the explanation, however, could be accepted by the users without their being able to perceive this failure.
Why does it happen?
First, it can occur that ML models output confident yet incorrect predictions. Such ML models can be said dysfunctional, raising the question of the relevance or potential misleadingness of generating explanations. Second, the XAI system is deemed dysfunctional when it fails to meet the mathematical objectives it has been designed to satisfy. This is related to the notion of unfaithful explanations i.e. that fail to adequately account for the behavior the ML model they are associated to, see e.g.  (Jacovi and Goldberg 2020; Li et al. 2023).
Many explanations are not generated through the minimization of a cost function, but are instead defined as closed-form formulas. Examples include saliency maps in Computer Vision (Selvaraju et al. 2017), influence functions (Koh and Liang 2017), partial dependence plots (Friedman 2001; Goldstein et al. 2015), or formal explanations (Darwiche and Hirth 2020; Audemard, Koriche, and Marquis 2020; Marques-Silva 2024). Such explanations can be considered as functional by design, and faithful.
On the other hand, some explanation methods rely on optimizing cost functions, see e.g., counterfactual example generation Wachter, Mittelstadt, and Russell (2018); Mothilal, Sharma, and Tan (2020); Laugel et al. (2018a) or surrogate-based methods (Ribeiro, Singh, and Guestrin 2016), and as such do not guarantee that the associated desiderata is satisfied. Often, there is no guarantee that a satisfying solution to these problems exists, as discussed for instance by Ye and Durrett (2022) for prompt-based explanations for Large Language Models. As a result, the explanation may be unfaithful to the model it aims at explaining.
Why is it a problem?
When an incorrect prediction is returned, although explanations can be useful for model calibration (Ye and Durrett 2022), they may be seen as potentially harmful, especially for users with low levels of awareness (Papenmeier, Englebienne, and Seifert 2019). An unfaithful explanation is also obviously problematic, as a lack of fidelity to the model may induce the user either to reject the system (undertrust) or to place unwarranted trust in it (overtrust(Papenmeier, Englebienne, and Seifert 2019; Colin et al. 2022).
Solutions.
Despite its critical importance, this faithfulness is still often overlooked in practice, by both users and ML practitioners (Kaur et al. 2020). Even when there is a will to control for faithfulness, the precise definition and evaluation of this notion remains elusive (Jacovi and Goldberg 2020; Laugel et al. 2018b; Li et al. 2023) and often at odds with other desired criteria (Han et al. 2023; Agarwal, Tanneru, and Lakkaraju 2024) amplifying the challenge. Like previous works, we argue that assessing faithfulness is crucial, and that this evaluation should precede all other assessments of the explanation: firstly, contrary to other assessments, it is a purely technical task, allowing ML developers to conduct it independently of end-users. Secondly, it serves as a foundational step for identifying and addressing any other potential issue. In the rest of the paper, the ML system is therefore assumed to be functional, i.e. to provide accurate predictions and faithful explanations.
4.3 Competing Explanations
We propose to name explanations competing when several parts of the explanation are contradicting with one another, if for instance they consist of several counterfactual examples.
Why does it happen?
We identify two scenarios when this can happen. Often, explanations are composed of several components, interacting with each other in various ways: e.g., a feature attribution vector represents the contributions of each feature, word or pixel to the prediction. Yet, numerous works show that more complex effects such as interactions or correlations between features are often not taken into account  (Gosiewska and Biecek 2019; Tsang, Rambhatla, and Liu 2020; Mase, Owen, and Seiler 2019; Casalicchio, Molnar, and Bischl 2019; Hooker, Mentch, and Zhou 2021). A contradiction may thus appear between the semantic relationship of two notions, their actual correlation in the data used to train the model, and the explanation returned by the system. The second scenario is when XAI systems generate, rather than a single explanation, a set of explanations, e.g. of counterfactual examples (Mothilal, Sharma, and Tan 2020) to provide richer insights to the user. In the counterfactual case, we connect the underlying notion of explanation diversity (see e.g. Laugel et al. (2023)) to the risk of getting competing explanations: these diverse explanations generally aim at suggesting the user various alternatives, i.e. a choice between several possible actions, but they may appear contradictory.
Why is it a problem?
Contradictions between two pieces of information may be perceived as confusing by the user, potentially leading them to reject the explanation (Suffian et al. 2022). For instance Bove et al. (2022) suggest competing explanations as one of the reasons for which users misunderstand the explanations returned by SHAP. Furthermore, in a non-cooperative setting (Bordt et al. 2022) where the objectives of the user and the ML developer are not aligned, this problem can also open up the risk of explanation manipulation through the selection of an explanation that is not in the best interest of the user, see e.g. Goethals, Martens, and Evgeniou (2023); Zhou and Joachims (2023) .
Solutions.
The proposed solutions can be grouped in two categories: the first one focuses on better XAI systems, either by adapting them to take into account correlations (Aas, Jullum, and Løland 2021; Salih et al. 2024), or enriching them, e.g. by computing, in addition to the usual feature importance vectors, feature interactions (Gosiewska and Biecek 2019; Jiang et al. 2025). A second type of enrichment consists in exploiting expert knowledge to contextualize feature contributions: informing the user, usually through the XUI, may allow to rationalize some confusion that can be caused by competing explanations (Bove et al. 2022).
4.4 Unstable Explanations
We call explanations unstable when there is an inconsistency, i.e. a contradiction, between explanations within a supposedly stable scenario: for instance, when producing local explanations for similar instances with similar outcomes, one might expect that the explanations should be similar as well (Jacovi and Goldberg 2020). Such explanation inconsistencies have been widely observed  (Alvarez-Melis and Jaakkola 2018a; Laugel et al. 2019; Yeh et al. 2019). Similarly to competing explanations unstable explanations are perceived originally as a technical failure of the explainer. Yet, as we describe below, it can also originate from the user, or from a combination of both the system and the user.
Why does it happen?
Unstable explanations are often considered as a technical failure of the explainer, viewed as a lack of robustness that needs to be fixed (Alvarez-Melis and Jaakkola 2018b; Slack et al. 2021b; Mishra et al. 2021; Kindermans et al. 2019). However, these inconsistencies can also be ascribed to the model to be explained (Dombrowski et al. 2019; Ghorbani, Abid, and Zou 2019; Alvarez-Melis and Jaakkola 2018b): the local behavior of the latter may indeed vary abruptly, due to the complexity of the task being modeled, the complexity of the model itself or its lack of robustness. Faithful explanations then reflect these steep changes, leading to an apparent lack of stability. Moreover, this issue of instability may conflate with user perception of the similarity between instances and the explanations they expect as a result. This similarity, that depends on the user knowledge and possible biases, may differ from the similarity considered by the ML system. Pushing the expectation of explanation stability to its extreme, explanations generated for identical observations are anticipated to be identical. However, post-hoc model-agnostic methods, be they local or global (e.g.  Ribeiro, Singh, and Guestrin (2016); Lundberg and Lee (2017); Wachter, Mittelstadt, and Russell (2018); Altmann et al. (2010)), often rely on a stochastic data generation step (Zhou, Hooker, and Wang 2021; Visani et al. 2022) that may cause instability. This comes in addition to the Roshomon effect, i.e. that several equally good but potentially drastically different solutions can coexist and therefore be selected as explanations (Hancox-Li 2020).
Why is it a problem?
Considering that faithful explanations reflect the state of ML model, depending on its potential causes discussed above, a lack of stability can either be seen as an actual failure or as a desired characteristic: for instance more local explanations are expected to be less stable (Molnar 2020; Yeh et al. 2019), with locality being a commonly expressed desideratum for explanations (Radensky et al. 2022). Thus, as for competing explanations, interpreting the lack of stability as a failure depends on the users needs, their knowledge of the explainer and their perception of how similar the explanations should be: a user not being aware of the locality-stability trade-off may see it as problematic but may not otherwise (Hancox-Li 2020). Still, instability may result in the user rejecting the explanation, or even the whole AI system, seeing it as a proof of unfairness by the model (Sharma, Henderson, and Ghosh 2020; Hickey, Di Stefano, and Vasileiou 2021), that may be abused by the organization providing the explanation (Goethals, Martens, and Evgeniou 2023; Zhou and Joachims 2023).
Solutions.
Various approaches have been proposed to address instability, depending on its source. To fix the stochastic instability of model-agnostic post-hoc explainers, most contributions focus on algorithmic modifications of the random data generation step they rely on, e.g. replacing it with a deterministic one (Zhou, Hooker, and Wang 2021; Zafar and Khan 2019), or through a reweighting strategy (Shankaranarayana and Runje 2019; Yeh et al. 2019). When instability originates from a misalignment between the user perception and the ML system representation, several works propose strategies to constrain the ML model during its training phase (Alvarez-Melis and Jaakkola 2018b; Dombrowski et al. 2019). On a different note, rather than mitigating the problem, several works propose to measure the explanation stability, arguing that it describes a notion of uncertainty that can be helpful for the user to better understand and use them (Gosiewska and Biecek 2019; Shankaranarayana and Runje 2019; Slack et al. 2021b; Visani et al. 2022).
4.5 Incompatible Explanations: the ”Disagreement Problem” in XAI
A third contradiction case occurs when several explainers are used in the same setting. We call them incompatible explanations, they correspond to the prevalent (Krishna et al. 2022) and well known Disagreement Problem, see e.g. Sundararajan and Najmi (2020); Han, Srinivas, and Lakkaraju (2022); Bordt et al. (2022); Neely et al. (2021).
Why does it happen?
Assuming that the explanations are faithful and stable, the most high-level root cause of this issue comes from the fact that the task of providing explanations, in particular in the post-hoc setting, is essentially underdetermined (Bordt et al. 2022): as mentioned in the introduction, the concept of ”good” explanation is elusive and has been formalized in numerous different ways, taking into account different types of desiderata. For instance, because they rely on different assumptions, LIME and SHAP explanations are expected to differ, even if both are faithful (Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022). Similar discussions apply to global feature attribution methods (Okeson et al. 2021) or counterfactual explanations (Goethals, Martens, and Evgeniou 2023). Going further, some differences between explanations can be attributed to discrepancies in the implementations of the supposedly same mathematical explanation objective, see e.g. Sundararajan and Najmi (2020) for the case of Shapley value-based explanations. Finally, a source of incompatibility can be attributed for the explanations generated using the same method, but different parameters. Indeed, XAI methods generally rely on hyperparameters that may not always be understood (if known at all) by the user, albeit heavily impacting the obtained explanations (Garreau and Luxburg 2020; Bansal, Agarwal, and Nguyen 2020). As a result, explanations can differ on multiple bases, for instance, in the case of feature score explanation, ranging from the top features being different to differences in order of importance or direction (Krishna et al. 2022).
Why (and when) is it a problem?
This disagreement between explanations is generally viewed as a problem (Sundararajan and Najmi 2020; Garreau and Luxburg 2020; Poyiadzi et al. 2021; Swamy et al. 2022; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Roy et al. 2022; Goethals, Martens, and Evgeniou 2023). Several user studies have noted it to be a source of confusion for users (Okeson et al. 2021; Krishna et al. 2022), resulting in their lowering trust in the system and therefore possibly leading to a reject of the system as a whole or a questionable selection of the proposed explanations, e.g. based on method popularity (Kaur et al. 2020). As for competing explanations, in a non-cooperative setting, this incompatibility between explanations may be leveraged by a malicious AI practitioner to rationalize unfair decisions by choosing the explanation most aligned to their objectives  (Slack et al. 2020; Aïvodji et al. 2019; Goethals, Martens, and Evgeniou 2023).
On the other hand, similarly to the unstable and competing cases, incompatible explanations, if faithful, can be viewed as an opportunity that may be leveraged for a better interaction with the system in a collaborative context. Indeed, explanation disagreement can be seen as a source of diversity, as noted by Laugel et al. (2023); Goethals, Martens, and Evgeniou (2023), which is viewed positively and helps understanding, as empirically shown by Bove et al. (2022) when combining counterfactual with global feature attributions. Other works have also leveraged explanation disagreement to reduce user overreliance to the model (Reingold, Shen, and Talati 2024). In other cases, disagreements between feature importance explanations are seen as an evidence of a lack of robustness on the side of the classifier, therefore used as a starting point for model auditing (Okeson et al. 2021) to improve its performance.
Solutions.
Intuitively, circumventing the issue of incompatibility may be simply done by hiding disagreements (Roy et al. 2022). On the contrary, other works propose to emphasize them through interfaces (Bove et al. 2023), sometimes arguing, like for stability, that the level of disagreement may be used as a measure of uncertainty to validate parts of the explanations (Krishna et al. 2022). Using the same idea, other works propose to aggregate various explanation methods (Bhatt and Moura 2021; Pirie et al. 2023; Decker et al. 2024) to provide more robust explanations, or even propose to train new models that minimize this disagreement (Schwarzschild et al. 2023).
5 Proposed Typology: User-specific
XAI Failures
This section discusses explanation failures that can occur when users misinterpret the explanations provided by a ML system with no technical failures: regardless of their quality, these explanations have been shown to sometimes fail in their explanatory objective (Cheng et al. 2019; Wang et al. 2019). We propose here a typology of failures that originates from inconsistent users’ inferences, distinguishing between three categories graphically represented in Fig. 2 and discussed in turn below: mismatch failures when there is a contradiction between the ML explanations and the users’ expectations in term of format; counterintuitive failures when this contradiction concerns the explanation content; and biased inferences failures when cognitive biases inherent to each user interfere with the explanations. We believe that these user-specific failures should be known so XAI designers can understand the users’ mental model processes and support their interpretation of the provided explanations.
Figure 2: Types of user-specific explanation failures in ML explanations. The ML model is supposed to give an accurate prediction and the explainer to generate faithful explanations. The cross indicates the step at which the failure takes place.
5.1 Mismatch
First, we propose to define mismatched explanations when the format of the extracted information does not meet the users’ expectations towards the ML system, leading to the remark ”This is not what I want” in Fig. 2.
Why does it happen?
Depending on the context of the interaction and the nature of the decision model’s outputs, users have been shown to have different needs and questions regarding the ML system, that may also vary depending on the model’s output  (Liao, Gruen, and Miller 2020; van der Waa et al. 2021). In parallel, there is a huge diversity in the forms explanations can take (see e.g. Dwivedi et al. (2023)), but explanation techniques do not necessarily take into account the context in which the explanations is sought by a user (Matarese, Rea, and Sciutti 2021). Therefore, mismatches between the user questions and the generated explanations can occur, on the explanation type (e.g., feature importance or rules), locality (e.g., local or global), goal (e.g., factual or causal) and on the information complexity (e.g., expressed for ML practitioners or lay users), to name a few.
Such mismatches reflect the lack of user-centricity in the conception of the XAI system. Actually, most XAI approaches are designed without evaluating whether the explanations satisfy the needs of real users (Doshi-Velez and Kim 2017; Keane et al. 2021). Instead, other criteria are used to evaluate the relevance of an XAI approach, such as the visual aspect of the explanations, its popularity or the ease of implementation (Mohseni, Zarei, and Ragan 2018; Barredo Arrieta et al. 2020). Such evaluations can bring to light that they fail to meet these needs: for example, during the co-design workshop for an AI-based diagnosis tool (Wang et al. 2019), many interviewed doctors reported that they would prefer alternative hypotheses (e.g., counterfactual examples) rather than factual explanations (e.g., local feature importance scores) that were initially implemented in the system.
Why is it a problem?
Mismatched explanations can lead to dissatisfaction and rejection of the whole ML system: it has been demonstrated that differences between initial expectations and actual experiences can affect both the user satisfaction and acceptance of a system (Bhattacherjee 2001). Along the same lines, the complexity of some explanation types can hinder users to understand the explanations and thus the adoption of the whole ML system (see e.g. Kaur et al. (2020) in the case of SHAP and GAM). It has also been demonstrated that users assign artificial agents human-like traits, and hence expect these agents to provide explanations using the same conceptual framework they are used to (De Graaf and Malle 2017). Yet, XAI solutions have not reached such a complete human-centered approach (Miller 2019; Matarese, Rea, and Sciutti 2021).
Solutions.
Mitigating the mismatched explanations somehow obviously relies on applying user-centered approaches when designing XAI systems: two examples include adjusting the types of explanations according to users’ expectations to improve their satisfaction and acceptance (Riveiro and Thill 2022; Pazzani et al. 2022) or integrating a lexical alignment step to improve the understanding of explanations provided by a conversational agent (Srivastava, Theune, and Catala 2023). Concept-bottleneck models (see e.g. Poeta et al. (2023); Zarlenga et al. (2024)) may be seen as examples of this family of approaches, when checking that the used concepts are meaningful for the users. Another type of solution aims at integrating social and cognitive science’s theories in a theoretical framework for XAI system, e.g. to provide personalized and contextualized explanations (Matarese, Rea, and Sciutti 2021; Byrne 2023).
5.2 Counterintuitive Explanations
We then propose to identify counterintuitive explanations when the explanation provided by the ML system is in contradiction with the prior knowledge of expert users (i.e., AI practitioners and domain experts). Contradiction here occurs at a content level, as opposed to the format level discussed in the case of counterintuitive failures in the previous section.
Why does it happen?
Prior knowledge may take various definitions (Dochy and Alexander 1995), we view it here as ”stored knowledge about the world that have been acquired by an individual” (Brod, Werkle-Bergner, and Shing 2013), including domain expertise and past experience. In an XAI context, users may find faithful explanations contradicting with such prior knowledge, leading them to perceive these explanations as different from what they anticipated (”This is not what I expected” in Fig. 2). This is all the more likely to happen as explanations are especially requested when the model’s output is perceived by users as abnormal or absurd (Riveiro and Thill 2022). In the same vein, past experience has been shown to lead to disagreement with the explanations (Sohn et al. 2019; Suffian et al. 2022).
Why is it a problem?
Counterintuitive explanations do not necessarily represent a failure: explanations that do not match user expectations can indeed be used to fix potential issues within the ML system (Kaur et al. 2020) or for knowledge discovery (Jiménez-Luna, Grisoni, and Schneider 2020). In other situations however, perceiving the provided explanations as counterintuitive can lead users to question the reliability of the prediction even when it is accurate, see e.g. Collaris, Vink, and van Wijk (2018); Palaniyappan Velumani et al. (2022) for studies in applied contexts. The ML system as a whole may be perceived more negatively (Cabitza et al. 2024; Nourani et al. 2021; Ebermann, Selisky, and Weibelzahl 2023), potentially impacting users’ willingness to engage with AI (Ebermann, Selisky, and Weibelzahl 2023). This aligns with findings from works in social sciences, which have shown that people tend to ignore information inconsistent with their beliefs from past experiences (Thagard 1989; Nickerson 1998).
Solutions.
Mitigating these failures requires to better align ML explanations and users’ prior knowledge. Some works in XAI thus argue for more personalized explanations (Ebermann, Selisky, and Weibelzahl 2023; Conati et al. 2021) that would directly integrate in their generation the user knowledge, e.g. expressed as features importance scores (Jeyasothy et al. 2022) or diagrams describing the reasoning process (Lim et al. 2025). Conversely, the predictive model itself can be changed so the generation of explanations that are aligned with users’ prior knowledge is facilitated (Rieger et al. 2020; Koh et al. 2020; Ross, Hughes, and Doshi-Velez 2017). In a different perspective, other works propose to co-design explanation interfaces together with experts so as to integrate both their needs and knowledge (Wang et al. 2019; Weitz et al. 2024).
5.3 Biased Inferences
Finally, we propose to identify explanation failures relying on biased inferences when users make inaccurate interpretations of the explanations, due to cognitive biases. As compared to counterintuitive failures that occur for expert users, for biased inferences we consider mainly lay users who do not have expert knowledge nor past experiences. Yet, such failures can occur for any type of users as cognitive biases are inherent to all humans.
Why does it happen?
Similarly to prior beliefs, biases can influence how users respond to different styles of explanations (Liao, Gruen, and Miller 2020; Miller 2019), and several cognitive biases have been shown to trigger inaccurate interpretations of explanations (Bertrand et al. 2022; Nourani et al. 2021; Wang et al. 2019). We discuss below some common cognitive biases with their consequences on the interpretation of explanations.
Why is it a problem?
First, some biases can trigger over reliance in ML explanations. It has been shown that having an explanation, regardless of its quality, increases trust (Hoff and Bashir 2015; Eiband et al. 2019; Lai and Tan 2019). In other examples, it is shown that longer, richer, explanations are found to be more plausible than shorter ones (Fürnkranz, Kliegr, and Paulheim 2020; He, Aishwarya, and Gadiraju 2025), and that users may believe they understand better than what they actually do (Rozenblit and Keil 2002; Mueller et al. 2019; Chromik et al. 2021; He, Kuiper, and Gadiraju 2023). On the other hand, other biases can trigger under reliance. For instance, the ”negativity bias” can cause lay users, in particular, to pay more attention or overweight negative information over positive one of the same strength (Kliegr, Bahník, and Fürnkranz 2021). It may lead users to pay more attention to negative outcomes of the ML system, thus eroding their trust (Pratto and John 1991). It has been demonstrated that showing the weaknesses of the system (e.g., competing explanations) or negative outcomes (e.g., a malignant diagnosis) early on can have a major influence on trust (Nourani et al. 2021). Finally, some other biases can trigger users to misapply the explanations: e.g., the ”insensitivity to sample size” bias may lead lay users to ignore the statistical significance of a statement (Fürnkranz, Kliegr, and Paulheim 2020); the ”availability” may lead lay users to believe that examples and events that easily come to mind are more representative than is actually the case (Wang et al. 2019); e.g., the ”primacy effect” bias may lead them to form an opinion based solely on the first piece of information received (Nourani et al. 2021).
Solutions.
Before mitigating these biased inferences, identifying them and measuring their effects on the users’ interpretation is a challenging task that can e.g. rely on comparing users’ objective and self-reported understanding (Cheng et al. 2019; Bove et al. 2022). Most approaches then rely on the design of appropriate XUI (Wang et al. 2019), for instance controlling what types of predictions users first see when interacting with the system, to mitigate the negative bias (Nourani et al. 2021).
6 Discussion
The typology of explanation failures presented in the previous section allows to understand why failures happen, how to mitigate them, and how to distinguish them from one another. In this section, we leverage this typology to discuss some key issues, and identify promising research directions.
6.1 Towards a Holistic XAI Approach
Observation: Some failures result from the interaction between components of the explanation process.
Many of the explanation failures discussed in the previous sections can be diagnosed as stemming from one of these components (model, explainer and user): they can for instance be due to a technical problem with the system. However, our analysis in the previous section also underlines that some errors actually arise from the interplay between the different components considered, rendering them incompatible: we discussed the interaction between the explainer and the user in Section 5 and discussed in Section 4 some cases of problematic interaction between the model and the explainer. For instance, issues like instability may highlight a mismatch between the decision model’s behavior (volatile, local), the implicit assumptions made during explanation design (stable boundaries), or the similarity between instances perceived by the user.
Consequence: subpar technical solutions.
In the case of these interactions between components, the failure is not caused by a deficiency in either component, but rather from their misalignment. Consequently, some technical solutions suggest adapting or replacing one of the components afterwards, sometimes taking a paradoxical turn: instead of questioning the explainer when observing a failure, it is sometime envisaged instead to train a new AI decision model that would be more adapted to this explainer. As an illustrative example, it has been proposed to build models with more stable behavior to mitigate instability issues of activation-based explainer (Alvarez-Melis and Jaakkola 2018b), or models constrained to minimize the disagreement between LIME and SHAP (Schwarzschild et al. 2023). This may seem surprising, as explainers are usually leveraged to generate insights about the model, not the other way round.
Going forward: towards a holistic design of the explanation process.
This problem underscores a significant challenge: the interplay between the elements of the system should be taken into account from the system’s inception. The three components of the system should be regarded as interconnected, rather than designed independently. This does not mean abandoning post-hoc methods but rather anticipating their integration in the overall explanation process. Multiple calls for a user-centric approach of XAI have been made, advocating for integrating user needs from the inception of the system and guiding the design of XAI methods (Wang et al. 2019; Ribera and Lapedriza 2019; Vellido 2020; Schmude et al. 2023). Nevertheless, more efforts should be pursued on the interaction between AI models and explainers. One possibility is to draw inspiration from research in Integrative Design for software systems (Tumer and Smidts 2010), proposing holistic design strategies for software systems. Design and monitoring of AI systems should thus be conceived in a holistic way, with any choice or change in the decision model prompting a reassessment of its compatibility with the explainer, and vice versa; and similarly for changes in user requirements.
6.2 Towards More Transparency and Personalization in XAI Systems
Observation: some failures happen because explainers are black-boxes, and both ML practitioners and users ignore their limitations.
The previous argument about the need to adopt a holistic approach of AI systems design also raises the question of the relevance of one-size-fits-all XAI solutions. The most well-known XAI methods (e.g. SHAP and LIME) are often conceived under assumptions of data-, model-, and user-agnosticity (i.e. absence of user focus), generally with the aim of providing more flexibility in their use. Yet, some of the system-specific failures discussed (see Sect. 4) and the previous argument about failures resulting from the interaction of several components suggest that these methods may not, in fact, be one-size-fits-all solutions. This adds up to the existing research questioning this user-agnosticity, suggesting that some types of explanations may not be suited to all user profiles (Hoff and Bashir 2015; Wang et al. 2019), or to all decision models (Alvarez-Melis and Jaakkola 2018b; Molnar et al. 2020) This restates that explainers face various limitations and various assumptions, that should be known and understood for proper use.
Consequence: ML practitioners and users misuse XAI systems.
Unfortunately, these limitations and assumptions of XAI systems are rarely known to users and even ML developers, as analyzed in the proposed typology. While this may seem intuitive for user-specific failures (e.g. overtrust issues in biased inferences), it also holds for system-based failures: for instance, some competing issues are caused by the user not knowing how the explainer handles correlations and interactions between features; some instability issues by their not understanding the tradeoff between stability and locality, nor what level of locality they wish; some incompatibility issues by their not understanding the differences actually captured by two explainers. Consequently, users may reject or misuse the explanation, not because it does not meet their needs, but because they do not understand how to use it, as shown by  Kaur et al. (2020).
Going forward: more ”transparency” in XAI methods.
Technical design choices of XAI systems and the assumptions they rely on heavily impact how the explanation should be interpreted. Overall, this further confirms the need for more effort in communicating on the core capabilities of XAI systems, e.g. through design principles such as ML transparency (Bove et al. 2022), and more generally improving algorithm literacy of non-ML users (Cabitza, Rasoini, and Gensini 2017; Chiang and Yin 2022; He, Aishwarya, and Gadiraju 2025). This also aligns with prior calls to further encourage interdisciplinary works for designing more transparent XAI systems, instead of building black-box explainers. One possible direction to pursue is improving the standardization of XAI methods (Haque, Islam, and Mikalef 2023), e.g. through the description of explanation methods in a factual way, akin to Model Cards (Mitchell et al. 2019).
6.3 Towards More XUI
Observation: some failures happen because users are limited in their interaction with explanations.
From a cognitive point of view, the explanation process is argued to be interactive (Miller 2019) and yet, at best, it is merely sequential in XAI. Most approaches are limited to the generation of factual information about the model’s behaviour or the predicted outcome (e.g., feature importance scores, counterfactual examples, etc.) and users are often not able to interact with it.
Consequence: external information and processes interfere with the explanations.
This causes the explanations to be potentially perceived by users as incomplete or incorrect (e.g., lack of knowledge or lack of transparency on the XAI method). As discussed in Sect. 5.3, users may thus draw on external information and cognitive processes to interpret the explanations, potentially leading to XAI failures. For instance, users may expect that there can be intuitive changes in counterfactual examples because they have experienced the same logical path in real life (Suffian et al. 2022).
Going forward: more user interfaces for explanations.
We believe that XUIs allow to better organize, complete and display ML explanations according to the user needs (Chromik and Butz 2021). Moreover, accounting for the users is key to design such interfaces, which thus forces AI practitioners to adopt a user-centered approach when conceiving ML systems (see Sect. 6.1). Previous studies have demonstrated the usefulness of visual interfaces to present ML explanations (see e.g., Szymanski, Millecamp, and Verbert (2021); Ooge, Kato, and Verbert (2022)) but there are other interaction modalities, in particular the conversational mode. As users have progressively become more familiar with conversational AIs (e.g. ChatGPT), the provided explanations may be presented in conversational interfaces and become a dialog, i.e. questions from the users and corresponding answers from the machine (Miller 2019; Ribera and Lapedriza 2019). This could open new perspectives for design principles such as introducing a narrative logic that allows temporizing the current information. For instance primary information can be controlled to mitigate the negativity bias (Nourani et al. 2021). We believe that studying such a modality for the display of explanations would allow to better understand users’ processes for analyzing and understanding ML explanations.
7 Conclusion
In this work, we have proposed a typology of XAI failures allowing to understand why failures happen, how to mitigate them, and how to distinguish them from one another. We believe it can help AI developers and designers better understand XAI systems and their limitations. Leveraging this typology, we have identified promising research avenues for XAI. In addition to these directions, future works will include investigating the potential interaction between multiple failures. Besides better understanding their connections to one another, the co-occurrence or superposition of several failures raises crucial questions regarding their consequences on user understanding. Furthermore, a better understanding of each failure and these interactions could be leveraged to formally propose a diagnostic framework to help identifying their origins in the explanation process.